In the reaction Mg(OH)²+2HCl--->MgCl²+2H²0 you were given 15 grams of Mg(OH)² and 25 grams of HCl.Which of the two reactants is the limiting reagents? How much MgCl² will be produced in the reaction?




1
Expert's answer
2022-04-21T15:18:04-0400

Solution:

The molar mass of Mg(OH)2 is 58.32 g/mol

The molar mass of HCl is 36.458 g/mol


Calculate moles of each reagent:

Moles of Mg(OH)2 = [15 g Mg(OH)2] × [1 mol Mg(OH)2 / 58.32 g Mg(OH)2] = 0.2572 mol Mg(OH)2

Moles of HCl = [25 g HCl] × [1 mol HCl / 36.458 g HCl] = 0.6857 mol HCl


Balanced chemical equation:

Mg(OH)2 + 2HCl → MgCl2 + 2H2O

According to stoichiometry:

1 mol of Mg(OH)2 reacts with 2 mol of HCl

Thus, 0.2572 mol of Mg(OH)2 react with:

[0.2572 mol Mg(OH)2] × [2 mol HCl / 1 mol Mg(OH)2] = 0.5144 mol HCl

However, initially there is 0.6857 mol of HCl (according to the task)

Therefore, Mg(OH)2 acts as limiting reagent and HCl is excess reagent


According to stoichiometry:

1 mol of Mg(OH)2 produces 1 mol of MgCl2

Thus, 0.2572 mol of Mg(OH)2 produce:

[0.2572 mol Mg(OH)2] × [1 mol MgCl2 / 1 mol Mg(OH)2] = 0.2572 mol MgCl2


The molar mass of MgCl2 is 95.211 g/mol

Therefore,

Mass of MgCl2 = [0.2572 mol MgCl2] × [95.211 g MgCl2 / 1 mol MgCl2] = 24.488 g MgCl2 = 24.5 g MgCl2

Mass of MgCl2 = 24.5 g


Answer:

The limiting reagent is Mg(OH)2

24.5 grams of MgCl2 will be produced in the reaction